# Jarník-type theorem for self-similar sets

### v1

@@ 6. Some Examples
The upper and lower bounds given in Lemma 6.3 are obtained precisely by estimating the $\mu$-measures of these two outer sets, as respectively given in [20, Theorem 9.1 and Lemma 12.7].
Motivated by this result together with Corollary 6.2, Chen [10] derived the following estimate for $\mu_{\bm{\omega}}(A_{Q}(\eta,\bf 0))$. Although Chen’s formulation is somewhat different from ours, it essentially implies the same result.
Proposition 6.4 ([10, Proposition 2.8]).
Let $K$ be a strongly irreducible self-simlar set of an IFS $\mathcal{F}$ satisfying the OSC and let $\mu$ be the corresponding $\delta$-Ahlfors regular self-similar measure. Let $Q\geq 1$ be a large integer and $0<\eta\leq Q^{-1/d}$. Then, there exists a constant $C_{\mathcal{F}}$ depending only on $\mathcal{F}$ such that for any ${\bm{\omega}}\in\Lambda^{*}$,
$Q\eta^{d}-C_{\mathcal{F}}\,\rho_{\bm{\omega}}^{-1-\frac{d\kappa}{d+1}}(Q^{-1}\eta)^{\frac{d\kappa}{d+1}}\ll\mu_{\bm{\omega}}\big(A_{Q}(\eta,{\bf 0})\big)\ll Q\eta^{d}+C_{\mathcal{F}}\,\rho_{\bm{\omega}}^{-1-\frac{d\kappa}{d+1}}(Q^{-1}\eta)^{\frac{d\kappa}{d+1}},$
@@
$Q\eta^{d}\geq 2C_{\mathcal{F}}\,\rho_{\bm{\omega}}^{-1-\frac{d\kappa}{d+1}}(Q^{-1}\eta)^{\frac{d\kappa}{d+1}}\quad\Longleftrightarrow\quad\rho_{\bm{\omega}}\geq 2C_{\mathcal{F}}\,Q^{-1}\eta^{-(d-\frac{d\kappa}{d+1})/(1+\frac{d\kappa}{d+1})}.$
@@
(6.5) $\rho_{\bm{\omega}}\geq 2C_{\mathcal{F}}\,Q^{-1}Q^{\alpha(d-\frac{d\kappa}{d+1})/(1+\frac{d\kappa}{d+1})}=2C_{\mathcal{F}}\,Q^{-(1+\frac{d\alpha(d+1)\kappa}{d+1+d\kappa}-d\alpha)}.$
@@
Since $K$ satisfies the open set condition, any ball $B$ contains a branch $K_{{\bm{\omega}}}$ of $K$ with $|B|\asymp|K_{{\bm{\omega}}}|$, and is itself contained in a finite union of branches of $K$, namely
$B\subset\bigcup_{i=1}^{i_{B}}K_{{\bm{\omega}}_{i}},$
where $i_{B}\asymp 1$ and $|B|\asymp|K_{{\bm{\omega}}_{i}}|$ for all $i$. By the self-simlarity and the Ahlfors regularity of $\mu$, we have
$|B|\asymp|K_{\bm{\omega}}|\quad\Longrightarrow\quad\mu(B)\asymp\mu(K_{\bm{\omega}}).$
Therefore, the corollary follows if, for any ${\bm{\omega}}\in\Lambda^{*}$ with $\rho_{\bm{\omega}}\geq Q^{-\beta}$ and $Q^{-\alpha}\leq\eta\leq Q^{-1/d}$,
$\mu\big(K_{\bm{\omega}}\cap A_{Q}(\eta,{\bf 0})\big)\asymp\mu(K_{\bm{\omega}})\cdot Q\eta^{d}.$
This is indeed the case, since by Proposition 6.4 and (6.5),
@@
This is always possible, since as $\alpha\downarrow 1$ the left-hand side converges to $\frac{2\kappa_{0}}{2+\kappa_{0}}>0$. Suppose that $\kappa_{0}$ is small enough so that there are nontrivial self-similar sets $K$ satisfying $\kappa(K)>\kappa_{0}$. Let $0<\delta_{0}<1$ be such that $\delta_{0}>1-\beta_{0}/2$ and $\kappa(K)>\kappa_{0}$ whenever $\operatorname{\dim_{H}}K\geq\delta_{0}$. The second restriction is possible because, by Theorem 6.1, the parameter $\kappa(K)$ is non-decreasing as the Hausdorff dimension of $K$ increases.
@@
Let $\mu_{b,\mathcal{D}}$ be the natrual normalized Hausdorff measure supported on $K(b,\mathcal{D})$, which will be referred to as the missing digits measure.
@@
$\dim_{l^{1}}\mu:=\sup\left\{s\geq 0:\sum_{\bm{\xi}\in\mathbb{Z}^{d}:|\bm{\xi}|\leq R}|\widehat{\mu}(\bm{\xi})|\ll R^{d-s}\text{ for all $R>0$}\right\},$
@@
$A(q,\eta,{\bm{\theta}}):=\{\mathbf{x}\in\mathbb{R}^{d}:\|q\mathbf{x}-{\bm{\theta}}\|<\eta\}.$
In [31, §6.2], Yu proved that for some $\varepsilon>0$ (which can be chosen arbitrarily small but, once fixed, remains constant), as long as $Q$ is sufficiently large and
@@
For any ${\bm{\omega}}\in\mathcal{D}^{*}:=\bigcup_{k\geq 1}\mathcal{D}^{k}$, denote by $K_{\bm{\omega}}$ and $\mu_{\bm{\omega}}$ the branch of $K$ and $\mu$ associated to $\bm{\omega}$}, repectively (see (6.3) for precise definitions). To prove a local version of (6.6), observe that for any ${\bm{\omega}}\in\mathcal{D}^{*}$,
(6.7) $|\widehat{\mu_{\bm{\omega}}}(\xi)|=|\widehat{\mu}(b^{-|{\bm{\omega}}|}\xi)|,$
where $|{\bm{\omega}}|$ denotes the length of $\bm{\omega}$}. Consequently, it is shown in [31, Pages 40-41] that for any $1/2<s_{1}<\dim_{l^{1}}\mu$, and any $\alpha$ and $\beta$ satisfying $1<\alpha<\frac{s_{1}}{1-s_{1}}$ and $0<\beta<\frac{1-(1+\alpha)(1-s_{1})}{s_{1}}$, we have
(6.8) $\sum_{q=Q}^{2Q}\mu_{\bm{\omega}}\big(A(q,\eta,\theta)\big)\asymp Q\eta$
whenever $Q$ is large enough (depending on $\alpha$ and $\beta$ only), $\theta\in\mathbb{R}$, $Q^{-\alpha}\leq\eta\leq Q^{-1}$ and $|K_{\bm{\omega}}|\geq Q^{-\beta}$. Note that $\mu_{\bm{\omega}}=\frac{\mu|_{K_{\bm{\omega}}}}{\mu(K_{\bm{\omega}})}$ for all ${\bm{\omega}}\in\mathcal{D}^{*}$ (see (6.3)). It follows that
(6.9) $\sum_{q=Q}^{2Q}\mu\big(K_{\bm{\omega}}\cap A(q,\eta,\theta)\big)=\mu(K_{\bm{\omega}})\cdot\sum_{q=Q}^{2Q}\mu_{\bm{\omega}}\big(A(q,\eta,\theta)\big)\asymp\mu(K_{\bm{\omega}})\cdot Q\eta.$
@@
$\mu\big(K_{\bm{\omega}}\cap A_{Q}(\eta,\theta)\big)\asymp\mu(K_{\bm{\omega}})\cdot Q\eta$
because the sets $A(q,\eta,\theta)$ may have overlap. In [31, Pages 57-58], after a carefully analysis of the measure of the overlaps, Yu was able to establish the above result for $\theta=0$. For arbitrary $\theta\in\mathbb{R}$, although some of his ideas are applicable, there are still several nontrivial difficulties.
@@
Suppose that $\dim_{l^{1}}\mu>1/2$. For any $1/2<s_{1}<\dim_{l^{1}}\mu$, let $1<\alpha<\frac{s_{1}}{1-s_{1}}$ and $0<\beta<\frac{1-(1+\alpha)(1-s_{1})}{s_{1}}$. Then, for any $\theta\in\mathbb{R}$, $\mu$ satisfies the $(\alpha,\beta,\theta)$-local counting property.
@@
To prove the lemma, it is enough to show that
(6.10) $\mu\big(K_{\bm{\omega}}\cap A_{Q}(\eta,\theta)\big)\asymp\mu(K_{\bm{\omega}})\cdot Q\eta,$
whenever $Q$ is sufficiently large (depending only on $\alpha$ and $\beta$), $\theta\in\mathbb{R}$, $Q^{-\alpha}\leq\eta\leq Q^{-1}$, and $|K_{\bm{\omega}}|\geq Q^{-\beta}$.
@@
(6.11) $\Longrightarrow\quad\frac{\int_{\{f_{Q}<k\}}f_{Q}{\,\mathrm{d}}\mu_{\bm{\omega}}}{k}\leq\mu_{\bm{\omega}}\big(A_{Q}(\eta,\theta)\big).$
@@
(6.12) $\int_{\{f_{Q}\geq k\}}f_{Q}{\,\mathrm{d}}\mu_{\bm{\omega}}\ll\frac{Q\eta}{k}+Q^{\rho}\eta\log_{2}Q,$
@@
Now, we prove the claim, and in the course of the proof it will become clear how the parameters $\alpha,\beta,s_{1}$ in the lemma arise. Let $Q^{-\alpha}\leq\eta<Q^{-1}/4$. Once the lower bound in (6.10) is established for this case, the corresponding lower bound for $\eta$ in the range $Q^{-1}/4\leq\eta\leq Q^{-1}$ follows immediately. Let $x\in\mathbb{R}$. If $f_{Q}(x)=m\geq 2$, then there exist exactly distinct $Q\leq q_{1}<\cdots<q_{m}\leq 2Q$ such that
@@
$\int_{\{f_{Q}\in[m,2m)\}}f_{Q}{\,\mathrm{d}}\mu_{\bm{\omega}}$ $\leq 2m\int\mathbbm{1}_{\bigcup_{q=1}^{Q/(m-1)}A(q,2\eta/(m-1),0)}{\,\mathrm{d}}\mu_{\bm{\omega}}$
(6.15) $\leq 2m\int\sum_{q=1}^{Q/(m-1)}\mathbbm{1}_{A(q,2\eta/(m-1),0)}{\,\mathrm{d}}\mu_{\bm{\omega}}.$
@@
We can modify $g$ by replacing the charater functions with smooth functions just as in the proof of [31, Theorem 4.1], and we see that it is possible to show that
(6.17) $\int_{[0,1]}g{\,\mathrm{d}}\mu_{\bm{\omega}}\ll\int_{[0,1]}g(x){\,\mathrm{d}}x+\sum_{\xi\in\mathbb{Z}:0<|\xi|\leq Q/\eta}|\widehat{\mu_{\bm{\omega}}}(\xi)\widehat{g}(\xi)|.$
@@
where $d(\xi)$ denote the number of divisors of $\xi$. The proof of this inequality can be found in [31, page 16]. This together with (6.7) gives
$\begin{split}\sum_{\xi\in\mathbb{Z}:0<|\xi|\leq Q/\eta}|\widehat{\mu_{\bm{\omega}}}(\xi)\widehat{g}(\xi)|&\ll\sum_{\xi\in\mathbb{Z}:0<|\xi|\leq Q/\eta}|\widehat{\mu_{\bm{\omega}}}(\xi)|d(\xi)\eta/m\\ &=\eta/m\cdot\sum_{\xi\in\mathbb{Z}:0<|\xi|\leq Q/\eta}|\widehat{\mu}(b^{-|{\bm{\omega}}|}\xi)|d(\xi).\end{split}$
It then follows from an argument in [31, Page 40] that for each $1/2<s_{1}<\dim_{l^{1}}\mu,$
$\sum_{0<|\xi|\leq Q/\eta}|\widehat{\mu}(b^{-|{\bm{\omega}}|}\xi)|d(\xi)\ll b^{s_{1}|{\bm{\omega}}|}Q^{1-s_{1}}\eta^{-(1-s_{1})}.$
@@
(6.19) $1<\alpha<\frac{s_{1}}{1-s_{1}}\mathrm{\quad and\quad}0<\beta<\frac{1-(1+\alpha)(1-s_{1})}{s_{1}},$
@@
(6.20) $b^{s_{1}|{\bm{\omega}}|}Q^{1-s_{1}}\eta^{-(1-s_{1})}\leq Q^{\rho}\quad\text{with $\rho=s_{1}\beta+(1+\alpha)(1-s_{1})\in(0,1)$}$
whenever $Q$ is large enough (depending on $\alpha$ and $\beta$ only), $Q^{-\alpha}\leq\eta\leq Q^{-1}$ and $|K_{\bm{\omega}}|=b^{-|{\bm{\omega}}|}\geq Q^{-\beta}$. This means that
@@
$\int_{\{f_{Q}\geq k\}}f_{Q}\,{\,\mathrm{d}}\mu_{\bm{\omega}}=\sum_{j\geq 0:\,2^{j}k\leq Q}\int_{\{f_{Q}\in[2^{j}k,2^{j+1}k)\}}f_{Q}\,{\,\mathrm{d}}\mu_{\bm{\omega}}.$
@@
$\begin{split}\sum_{j\geq 0:\,2^{j}k\leq Q^{1-\rho^{\prime}}}\int_{\{f_{Q}\in[2^{j}k,2^{j+1}k)\}}f_{Q}\,{\,\mathrm{d}}\mu_{\bm{\omega}}&\ll\sum_{j\geq 0:\,2^{j}k\leq Q}\biggl(\frac{Q\eta}{2^{j}k}+Q^{\rho}\eta\biggr)\\ &\ll\frac{Q\eta}{k}+Q^{\rho}\eta\log_{2}Q,\end{split}$
@@
Using the upper bound $\beta<\frac{1-(1+\alpha)(1-s_{1})}{s_{1}}$, we obtain
$\begin{split}1-\alpha+(1+\alpha-\beta)\delta&\geq 1-\alpha+\bigg(1+\alpha-\frac{1-(1+\alpha)(1-s_{1})}{s_{1}}\bigg)\delta\\ &=1+\alpha(\delta/s_{1}-1).\end{split}$
The desired conclusion then follows from the fact that, for missing digits measures, $\delta\geq\dim_{l^{1}}\mu>s_{1}$ (see [31, Lemma 1.4 (1) and §3.6]).
@@
(2) Since $\operatorname{\dim_{H}}K\dim_{l^{1}}\mu=\delta\cdot\dim_{l^{1}}\mu>1/2$, we can choose $s_{1}$ such that $1/2<s_{1}<\dim_{l^{1}}\mu$ and $\delta s_{1}>1/2$. With this choice of $s_{1}$, we further choose parameters $\alpha$ and $\beta$ such that
(6.22) $1<\alpha<\frac{2\delta s_{1}-s_{1}}{1-s_{1}}\mathrm{\quad and\quad}2(1-\delta)\leq\beta<\frac{1-(1+\alpha)(1-s_{1})}{s_{1}}.$
In order for the set of admissible $\beta$ values to be nonempty, and thus for $\delta\geq 1-\beta/2$ to hold, the upper bound that $\alpha$ can take must be less than $\frac{s_{1}}{1-s_{1}}$. Indeed, the set of admissible $\beta$ values is nonempty if and only if
$2(1-\delta)<\frac{1-(1+\alpha)(1-s_{1})}{s_{1}},$
@@
$\alpha<\frac{2\delta s_{1}-s_{1}}{1-s_{1}},$

### v2

@@ 6. Some Examples
The upper and lower bounds given in Lemma 6.3 are obtained precisely by estimating the $\mu$-measures of these two outer sets, as respectively given in [25, Theorem 9.1 and Lemma 12.7].
Motivated by this result together with Corollary 6.2, Chen [12] derived the following estimate for $\mu_{\omega}(A_{Q}(\eta,\bf 0))$. Although Chen’s formulation is somewhat different from ours, it essentially implies the same result.
Proposition 6.4 ([12, Proposition 2.8]).
Let $K$ be a strongly irreducible self-simlar set of an IFS $\mathcal{F}$ satisfying the OSC and let $\mu$ be the corresponding $\delta$-Ahlfors regular self-similar measure. Let $Q\geq 1$ be a large integer and $0<\eta\leq Q^{-1/d}$. Then, there exists a constant $C_{\mathcal{F}}$ depending only on $\mathcal{F}$ such that for any $\omega\in\Lambda^{*}$,
$Q\eta^{d}-C_{\mathcal{F}}\,\rho_{\omega}^{-1-\frac{d\kappa}{d+1}}(Q^{-1}\eta)^{\frac{d\kappa}{d+1}}\ll\mu_{\omega}\big(A_{Q}(\eta,{\bf 0})\big)\ll Q\eta^{d}+C_{\mathcal{F}}\,\rho_{\omega}^{-1-\frac{d\kappa}{d+1}}(Q^{-1}\eta)^{\frac{d\kappa}{d+1}},$
@@
$Q\eta^{d}\geq 2C_{\mathcal{F}}\,\rho_{\omega}^{-1-\frac{d\kappa}{d+1}}(Q^{-1}\eta)^{\frac{d\kappa}{d+1}}\quad\Longleftrightarrow\quad\rho_{\omega}\geq 2C_{\mathcal{F}}\,Q^{-1}\eta^{-(d-\frac{d\kappa}{d+1})/(1+\frac{d\kappa}{d+1})}.$
@@
(6.5) $\rho_{\omega}\geq 2C_{\mathcal{F}}\,Q^{-1}Q^{\alpha(d-\frac{d\kappa}{d+1})/(1+\frac{d\kappa}{d+1})}=2C_{\mathcal{F}}\,Q^{-(1+\frac{d\alpha(d+1)\kappa}{d+1+d\kappa}-d\alpha)}.$
@@
Since $K$ satisfies the open set condition, any ball $B$ contains a branch $K_{\omega}$ of $K$ with $|B|\asymp|K_{\omega}|$, and is itself contained in a finite union of branches of $K$, namely
$B\subset\bigcup_{i=1}^{i_{B}}K_{\omega_{i}},$
where $i_{B}\asymp 1$ and $|B|\asymp|K_{\omega_{i}}|$ for all $i$. By the self-simlarity and the Ahlfors regularity of $\mu$, we have
$|B|\asymp|K_{\omega}|\quad\Longrightarrow\quad\mu(B)\asymp\mu(K_{\omega}).$
Therefore, the corollary follows if, for any $\omega\in\Lambda^{*}$ with $\rho_{\omega}\geq Q^{-\beta}$ and $Q^{-\alpha}\leq\eta\leq Q^{-1/d}$,
$\mu\big(K_{\omega}\cap A_{Q}(\eta,{\bf 0})\big)\asymp\mu(K_{\omega})\cdot Q\eta^{d}$
whenever $Q$ is sufficiently large (depending only on $\alpha$ and $\beta$). This is indeed the case, since by Proposition 6.4 and (6.5),
@@
This is always possible, since as $\alpha\downarrow 1$ the left-hand side converges to $\frac{2\kappa_{0}}{2+\kappa_{0}}>0$. Suppose that $\kappa_{0}>0$ is small enough so that there are nontrivial self-similar sets $K$ satisfying $\kappa(K)>\kappa_{0}$. Let $0<\delta_{0}<1$ be such that $\delta_{0}>1-\beta/2$ and $\kappa(K)>\kappa_{0}$ whenever $\operatorname{\dim_{H}}K\geq\delta_{0}$. The second restriction is possible because, by Theorem 6.1, the parameter $\kappa(K)$ is non-decreasing as the Hausdorff dimension of $K$ increases.
@@
Let $\mu_{b,\mathcal{D}}$ be the natural normalized Hausdorff measure supported on $K(b,\mathcal{D})$, which will be referred to as the missing digits measure.
@@
$\dim_{l^{1}}\mu:=\sup\left\{s\geq 0:\sum_{\bm{\xi}\in\mathbb{Z}^{d}:|\bm{\xi}|\leq M}|\widehat{\mu}(\bm{\xi})|\ll M^{d-s}\text{ for all $M>0$}\right\},$
@@
(6.6) $A(q,\eta,{\bm{\theta}}):=\{\mathbf{x}\in\mathbb{R}^{d}:\|q\mathbf{x}-{\bm{\theta}}\|<\eta\}.$
In [42, §6.2], Yu proved that for some $\varepsilon>0$ (which can be chosen arbitrarily small but, once fixed, remains constant), as long as $Q$ is sufficiently large and
@@
For any $\omega\in\mathcal{D}^{*}:=\bigcup_{k\geq 1}\mathcal{D}^{k}$, denote by $K_{\omega}$ and $\mu_{\omega}$ the branch of $K$ and $\mu$ associated to $\omega$, repectively (see (6.3) for precise definitions). To prove a local version of (6.7), observe that for any $\omega\in\mathcal{D}^{*}$,
(6.8) $|\widehat{\mu_{\omega}}(\xi)|=|\widehat{\mu}(b^{-|\omega|}\xi)|,$
where $|\omega|$ denotes the length of $\omega$. Consequently, it is shown in [42, Pages 40-41] that for any $1/2<\gamma<\dim_{l^{1}}\mu$, and any $\alpha$ and $\beta$ satisfying $1<\alpha<\frac{\gamma}{1-\gamma}$ and $0<\beta<\frac{1-(1+\alpha)(1-\gamma)}{\gamma}$, we have
(6.9) $\sum_{q=Q}^{2Q}\mu_{\omega}\big(A(q,\eta,\theta)\big)\asymp Q\eta$
whenever $Q$ is large enough (depending on $\alpha$ and $\beta$ only), $\theta\in\mathbb{R}$, $Q^{-\alpha}\leq\eta\leq Q^{-1}$ and $|K_{\omega}|\geq Q^{-\beta}$. Note that $\mu_{\omega}=\frac{\mu|_{K_{\omega}}}{\mu(K_{\omega})}$ for all $\omega\in\mathcal{D}^{*}$ (see (6.3)). It follows that
(6.10) $\sum_{q=Q}^{2Q}\mu\big(K_{\omega}\cap A(q,\eta,\theta)\big)=\mu(K_{\omega})\cdot\sum_{q=Q}^{2Q}\mu_{\omega}\big(A(q,\eta,\theta)\big)\asymp\mu(K_{\omega})\cdot Q\eta.$
@@
$\mu\big(K_{\omega}\cap A_{Q}(\eta,\theta)\big)\asymp\mu(K_{\omega})\cdot Q\eta$
because the sets $A(q,\eta,\theta)$ may have overlaps. In [42, Pages 57-58], after a carefully analysis of the measure of the overlaps, Yu was able to establish the above result for $\theta=0$. For arbitrary $\theta\in\mathbb{R}$, although some of his ideas are applicable, there are still several nontrivial difficulties.
@@
Suppose that $\dim_{l^{1}}\mu>1/2$. For any $1/2<\gamma<\dim_{l^{1}}\mu$, let $1<\alpha<\frac{\gamma}{1-\gamma}$ and $0<\beta<\frac{1-(1+\alpha)(1-\gamma)}{\gamma}$. Then, for any $\theta\in\mathbb{R}$, $\mu$ satisfies the $(\alpha,\beta,\theta)$-local counting property.
@@
Let $\theta\in\mathbb{R}$. To prove the lemma, it is enough to show that
(6.11) $\mu\big(K_{\omega}\cap A_{Q}(\eta,\theta)\big)\asymp\mu(K_{\omega})\cdot Q\eta,$
whenever $Q$ is sufficiently large (depending only on $\alpha$ and $\beta$), $Q^{-\alpha}\leq\eta\leq Q^{-1}$, and $|K_{\omega}|\geq Q^{-\beta}$.
@@
(6.12) $\Longrightarrow\quad\frac{\int_{\{f_{Q}<k\}}f_{Q}{\,\mathrm{d}}\mu_{\omega}}{k}\leq\mu_{\omega}\big(A_{Q}(\eta,\theta)\big).$
@@
(6.13) $\int_{\{f_{Q}\geq k\}}f_{Q}{\,\mathrm{d}}\mu_{\omega}\ll\frac{Q\eta}{k}+Q^{\rho}\eta\log_{2}Q,$
@@
Now, we prove the claim, and in the course of the proof it will become clear how the parameters $\alpha,\beta,\gamma$ in the lemma arise. Let $Q^{-\alpha}\leq\eta<Q^{-1}/4$. Once the lower bound in (6.11) is established for this case, the corresponding lower bound for $\eta$ in the range $Q^{-1}/4\leq\eta\leq Q^{-1}$ follows immediately. Let $x\in\mathbb{R}$. If $f_{Q}(x)=m\geq 2$, then there exist exactly distinct $Q\leq q_{1}<\cdots<q_{m}\leq 2Q$ such that
@@
$\int_{\{f_{Q}\in[m,2m)\}}f_{Q}{\,\mathrm{d}}\mu_{\omega}$ $\leq 2m\int\mathbbm{1}_{\bigcup_{q=1}^{Q/(m-1)}A(q,2\eta/(m-1),0)}{\,\mathrm{d}}\mu_{\omega}$
(6.16) $\leq 2m\int\sum_{q=1}^{Q/(m-1)}\mathbbm{1}_{A(q,2\eta/(m-1),0)}{\,\mathrm{d}}\mu_{\omega}.$
@@
We can modify $g$ by replacing the charater functions with smooth functions just as in the proof of [42, Theorem 4.1], and we see that it is possible to show that
(6.18) $\int_{[0,1]}g{\,\mathrm{d}}\mu_{\omega}\ll\int_{[0,1]}g(x){\,\mathrm{d}}x+\sum_{\xi\in\mathbb{Z}:0<|\xi|\leq Q/\eta}|\widehat{\mu_{\omega}}(\xi)\widehat{g}(\xi)|.$
@@
where $d(\xi)$ denote the number of divisors of $\xi$. The proof of this inequality can be found in [42, page 16]. This together with (6.8) gives
$\begin{split}\sum_{\xi\in\mathbb{Z}:0<|\xi|\leq Q/\eta}|\widehat{\mu_{\omega}}(\xi)\widehat{g}(\xi)|&\ll\sum_{\xi\in\mathbb{Z}:0<|\xi|\leq Q/\eta}|\widehat{\mu_{\omega}}(\xi)|d(\xi)\eta/m\\ &=\eta/m\cdot\sum_{\xi\in\mathbb{Z}:0<|\xi|\leq Q/\eta}|\widehat{\mu}(b^{-|\omega|}\xi)|d(\xi).\end{split}$
It then follows from an argument in [42, Page 40] that for each $1/2<\gamma<\dim_{l^{1}}\mu,$
(6.20) $\sum_{0<|\xi|\leq Q/\eta}|\widehat{\mu}(b^{-|\omega|}\xi)|d(\xi)\ll b^{\gamma|\omega|}Q^{1-\gamma}\eta^{-(1-\gamma)}.$
@@
(6.21) $1<\alpha<\frac{\gamma}{1-\gamma}\mathrm{\quad and\quad}0<\beta<\frac{1-(1+\alpha)(1-\gamma)}{\gamma},$
@@
(6.22) $b^{\gamma|\omega|}Q^{1-\gamma}\eta^{-(1-\gamma)}\leq Q^{\rho}\quad\text{with $\rho=\gamma\beta+(1+\alpha)(1-\gamma)\in(0,1)$}$
whenever $Q$ is large enough (depending on $\alpha$ and $\beta$ only), $Q^{-\alpha}\leq\eta\leq Q^{-1}$ and $|K_{\omega}|=b^{-|\omega|}\geq Q^{-\beta}$. This means that
@@
$\int_{\{f_{Q}\geq k\}}f_{Q}\,{\,\mathrm{d}}\mu_{\omega}=\sum_{j\geq 0:\,2^{j}k\leq Q}\int_{\{f_{Q}\in[2^{j}k,2^{j+1}k)\}}f_{Q}\,{\,\mathrm{d}}\mu_{\omega}.$
@@
$\begin{split}\sum_{j\geq 0:\,2^{j}k\leq Q^{1-\rho^{\prime}}}\int_{\{f_{Q}\in[2^{j}k,2^{j+1}k)\}}f_{Q}\,{\,\mathrm{d}}\mu_{\omega}&\ll\sum_{j\geq 0:\,2^{j}k\leq Q}\biggl(\frac{Q\eta}{2^{j}k}+Q^{\rho}\eta\biggr)\\ &\ll\frac{Q\eta}{k}+Q^{\rho}\eta\log_{2}Q,\end{split}$
@@
Using the upper bound $\beta<\frac{1-(1+\alpha)(1-\gamma)}{\gamma}$, we obtain
$\begin{split}1-\alpha+(1+\alpha-\beta)\delta&\geq 1-\alpha+\bigg(1+\alpha-\frac{1-(1+\alpha)(1-\gamma)}{\gamma}\bigg)\delta\\ &=1+\alpha(\delta/\gamma-1).\end{split}$
The desired conclusion then follows from the fact that, for missing digits measures, $\delta\geq\dim_{l^{1}}\mu>\gamma$ (see [42, Lemma 1.4 (1) and §3.6]).
@@
(2) Since $\operatorname{\dim_{H}}K\dim_{l^{1}}\mu=\delta\cdot\dim_{l^{1}}\mu>1/2$, we can choose $\gamma$ such that $1/2<\gamma<\dim_{l^{1}}\mu$ and $\delta\gamma>1/2$. With this choice of $\gamma$, we further choose parameters $\alpha$ and $\beta$ such that
(6.24) $1<\alpha<\frac{2\delta\gamma-\gamma}{1-\gamma}\mathrm{\quad and\quad}2(1-\delta)\leq\beta<\frac{1-(1+\alpha)(1-\gamma)}{\gamma}.$
In order for the set of admissible $\beta$ values to be nonempty, and thus for $\delta\geq 1-\beta/2$ to hold, the upper bound that $\alpha$ can take must be less than $\frac{\gamma}{1-\gamma}$. Indeed, the set of admissible $\beta$ values is nonempty if and only if
$2(1-\delta)<\frac{1-(1+\alpha)(1-\gamma)}{\gamma},$
@@
$\alpha<\frac{2\delta\gamma-\gamma}{1-\gamma},$